OpenStudy (anonymous):

The graph shows the functions f(x), p(x), and g(x): Graph of function g of x is y is equal to 1.8 to the power of x. The straight line f of x joins ordered pairs 6, 2 and 2, minus 2 and is extended on both sides. The straight line p of x joins the ordered pairs 3, 6 and 6, 2 and is extended on both sides. Part A: What is the solution to the pair of equations represented by p(x) and f(x)? (3 points) Part B: Write any two solutions for f(x). (3 points) Part C: What is the solution to the equation p(x) = g(x)? Justify your answer. (4 points)

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

i just seriously dont get any of this

OpenStudy (anonymous):

part a = f(x)= (6,2) and (2,-2) p(x)= (3,6) and( 6,2) this is all i have so far.

hero (hero):

So you are given \(g(x) = 1.8^x\) \(f(x)\) includes points (6,2) and (2,-2) \(p(x)\) includes points (3,6) and (6,2) What point does \(f(x)\) and \(p(x)\) have in common?

OpenStudy (anonymous):

the points (6,2)

hero (hero):

Yes, so basically, since \(f(x)\) and \(p(x)\) have (6,2) in common, it means that's the intersection point.

hero (hero):

The solution point.

OpenStudy (anonymous):

....

OpenStudy (anonymous):

this is all i got so far part a = f(x)= (6,2) and (2,-2) p(x)= (3,6) and( 6,2

hero (hero):

Let me post a graph of the situation.

OpenStudy (anonymous):

all i need for part a is What is the solution to the pair of equations represented by p(x) and f(x)?

OpenStudy (anonymous):

so what are the equations for f(x) and p(x) and wahts the solution to them. thats what i need for part a

hero (hero):

@alyssa_fields50, please do not panic.

OpenStudy (anonymous):

i dont need a graph

OpenStudy (anonymous):

i dont get any of this. i dont understand anything, im going to fail.

hero (hero):

Observe where the point (6,2) lies in relation to f(x) and p(x).

OpenStudy (anonymous):

i dont get it.

hero (hero):

The point (6,2) is the intersection point of f(x) and g(x). The intersection point of two functions is also the solution point

hero (hero):

Just think about it and try to understand it.

OpenStudy (anonymous):

i just need a solution to the pair of equations.

OpenStudy (anonymous):

im so lost. please try to keep helping me.

OpenStudy (anonymous):

is it 6,3?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

6,2?

hero (hero):

6,2 is the solution point as I alluded to earlier.

OpenStudy (anonymous):

part a = f(x)= (6,2) and (2,-2) p(x)= (3,6) and( 6,2) so the soltition to part a = (6,2) is this correct for part a?

hero (hero):

Yes, (6,2) is correct for part a

OpenStudy (anonymous):

so is this what youd put for part a part a = f(x)= (6,2) and (2,-2) p(x)= (3,6) and( 6,2) so the soltition to part a = (6,2)

hero (hero):

too many equal signs.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

what exactly would you put for part a

hero (hero):

Here's exactly what you write for part A: Q: What is the solution to the pair of equations represented by p(x) and f(x)? A: The solution to the pair of equations represented by p(x) and f(x) is (6,2).

OpenStudy (anonymous):

hmmm....

OpenStudy (anonymous):

can you help with the other parts?

OpenStudy (anonymous):

@Hero